The cost, C, of a manufacturing and selling x units of a product is C = 23x+73, and the corresponding revenue, R, is= x^2-35. Find the break-even value of x?
A. 4
B. 31
C. 31 and 4
D. 27

Answer by nerdybill(7384)

The cost, C, of a manufacturing and selling x units of a product is C = 23x+73, and the corresponding revenue, R, is= x^2-35. Find the break-even value of x?
.
Break-even point is when
C=R
23x+73 = x^2-35
73 = x^2-23x-35
0 = x^2-23x-108
0 = (x+4)(x-27)
x = {-4, 27}
Throw out the negative solution (extraneous) leaving:
x = 27 units